I think this problem is vague because it doesnt say whether or not the card is put back. I am assuming it is...but would appreciate your perspective on solving both ways.
"Your friend makes up a game with a single suit of cards in a deck (13 cards). He says every time you pick a random card, someone wins money. If an even card is picked, he will pay you $10. If a card that is NOT EVEN is picked, you pay him $5. Explain whether you will or will NOT play his game.

Answer by edjones(8007)

There are 5 even cards 2,4,6,8,10 and 8 non-even cards A,3,5,7,9,J,Q,K
5*10=50 you will win if you pick each even card once.
8*5=40 you will lose if you pick each non-even card once.
You have an expectation of winning $10 after 13 picks or $10/13=~ $0.77 a turn.
.
Ed